# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ F.3 Derivation of von Mises expansion in Theorem 1 and Theorem 2
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
Using that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $\mu_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$, we can show that
$E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
@@
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mathcal{T}_{P}(q_{P})(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
The first result then follows. In the case where $P,\widehat{P}\in\mathcal{P}_{H}$, we have that $\mu_{P}=\mathcal{T}_{P}(q_{P,H})$ and $\widehat{\mu}=\widehat{\mathcal{T}}(\widehat{q}_{H})$. In this case, the expression simplifies to:
@@
Appendix G Proofs for Section 3 on semiparametric DRL
G.1 Asymptotic linearity of DRL estimator under correct specification
In the following conditions and theorem, let $\varphi_{n,H}$ denote the estimator of the IF $\varphi_{0}$ from Theorem 1, obtained by plugging in our nuisance estimators.
@@ G.1 Asymptotic linearity of DRL estimator under correct specification
G.2 Asymptotic linearity and efficiency of model-robust DRL estimator
@@
$R_{n,H}^{*}(P_{0})$ $:=P_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{n,H})\right\}(\mu_{0}-\mu_{n})\right]$
$\quad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\}(\mathcal{T}_{0}(q_{0,H})-\mathcal{T}_{0}(q_{n,H}^{\pi}))\right]$
$\quad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\}(\mathcal{T}_{0}(q_{n,H}^{\pi})-\widehat{\mathcal{T}}_{n}(q_{n,H}^{\pi}))\right].$
@@
Appendix H Proofs for Section 5 on calibrated FQI
H.1 Lemma bounding approximation error of estimated features
In the following lemma, let $X$ be a covariate and $Y\in\mathbb{R}$ be an outcome. For feature transformations $\varphi_{n},\varphi_{0}$, denote $f_{(\varphi_{n},\varphi_{0})}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x),\varphi_{0}(X)=\varphi_{0}(x)]$, $f_{\varphi_{n}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x)]$, and $f_{\varphi_{0}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{0}(X)=\varphi_{0}(x)]$.
@@ H.1 Lemma bounding approximation error of estimated features
$\left|g(\varphi_{n}(x),\varphi_{0}(x))-E[g(\varphi_{n}(X),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$ $=\left|E[g(\varphi_{n}(x),\varphi_{0}(x))-g(\varphi_{n}(x),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$
@@
Now, for some function $f$, suppose that $(\varphi_{n}(x),\varphi_{0}(x))\mapsto(\Pi_{\varphi_{n},\varphi_{0}}f)(x)$ is Lipschitz continuous. Then, defining $g:(\widehat{m},m)\mapsto E_{0}[f(X)|\varphi_{n}(X)=\widehat{m},\varphi_{0}(X)=m,\mathcal{D}_{n}]$ and noting by the law of iterated expectation that $\Pi_{\varphi_{n}}\Pi_{\varphi_{n},\varphi_{0}}f=\Pi_{\varphi_{n}}f$, we obtain the following pointwise error bound:
@@
H.2 Proofs of main results
@@
We introduce the following notation. For any feature transformation $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, let $d_{0,\phi}=T_{\phi}(\alpha_{0,\phi})$, where $\alpha_{0,\phi}:=\operatorname*{argmin}_{\alpha\in\overline{H}_{\phi}}E_{0}\left[\{\mathcal{T}_{0,\phi}(f\circ\phi)(A_{0},S_{0})\}^{2}-2m(S_{0},A_{0},\alpha)\right]$ is the Riesz representer for the function class $H_{\phi}:=\{f\circ\phi;f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$ induced by $\phi$. Denote $\mu_{0,\phi}$ by the map $(a,s)\mapsto E_{0}[Y_{0}\mid\phi(A_{0},S_{0})=\phi(a,s)]$. Let $\widetilde{\alpha}_{0,q_{n}}:=\operatorname*{argmin}_{f\in\overline{H}_{q_{n}}}\left\|\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(f\circ q_{n})\right\|_{P_{0}}$. Let $\widetilde{d}_{0,q_{n}}$ be either $\mathcal{T}_{0,q_{n}}(\alpha_{0,q_{n}})$ or $\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}})$.
@@
$=\langle d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$\quad+\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
Note $\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})=\mu_{0}$ by correct specification of $H_{(q_{n},q_{0})}$ for $q_{0}$. Thus, by the law of total expectation, we have
$\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$ $=\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$=\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-\mu_{0}\rangle_{P_{0}}$
@@
where the final equality follows from the orthogonality conditions of the Bellman projection $q_{0,q_{n}}$ of $q_{0}$ with respect to the norm $\|\mathcal{T}_{0,q_{n}}(\cdot)\|_{P_{0}}$ and the fact that $\widetilde{d}_{0,q_{n}}$ is in the closed range of $\mathcal{T}_{0,q_{n}}$. Hence,
$\Psi_{n}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
@@
$|\Psi_{n}(P_{0})-\Psi(P_{0})|$ $\leq\|d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}}\|_{P_{0}}\cdot\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}}.$
For the second part of the theorem, take $\widetilde{d}_{0,q_{n}}$ equal to $\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}})$. By the invertibility condition in D1, $\mathcal{T}_{0,q_{n}}(\overline{H}_{q_{n}})$ equals $\overline{H}_{q_{n}}$ from the first part of (D2). Hence, we have that $\widetilde{d}_{0,q_{n}}$ equals the projection $\operatorname*{argmin}_{d\in\overline{H}_{q_{n}}}\|d_{0,(q_{n},q_{0})}-d\|_{P_{0}}.$ Next, using the second part of (D2), we apply Lemma 14 with $Y:=d_{0,(q_{n},q_{0})}$, $\varphi_{n}:=q_{n}$, and $\varphi_{0}:=q_{0}$ to conclude that
$\|d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}}\|_{P_{0}}\lesssim\|q_{n}-q_{0}\|_{P_{0}}.$
@@
$\sum_{i=1}^{n}f(q_{n}(A_{0,i},S_{0,i}))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}(A_{0,i},S_{0,i})\}=0.$
Taking $f$ such that $f\circ q_{n}=\widetilde{d}_{0,q_{n}}$, we find that
$\sum_{i=1}^{n}\widetilde{d}_{0,q_{n}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}(A_{0,i},S_{0,i})\}=0.$
Therefore, the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n})$ is equal to the DRL estimator:
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n})+\sum_{i=1}^{n}\widetilde{d}_{0,q_{n}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}(A_{0,i},S_{0,i})\}.$
@@
[…]
@@
[…]
@@
[…]
@@
$\quad+\langle{d}_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{n}^{*})\rangle_{P_{0}}$
@@
Appendix I Proofs for ADML in Section B
@@ I.1 Proof of Theorem 10
Observe that $\psi_{n,H_{n}}=\Psi_{H_{n}}(\widehat{P_{n}})+P_{n}\varphi_{n,H_{n}}$, where $\widehat{P}_{n}\in\mathcal{P}$ is any distribution such that $q_{\widehat{P}_{n},H_{n}}=q_{n,H_{n}}$, $\mu_{\widehat{P}_{n}}=\widehat{\mathcal{T}}_{n}(q_{n,H_{n}})$, and $T_{\widehat{P}_{n}}(\alpha_{\widehat{P}_{n},H_{n}})=T_{\widehat{P}_{n}}(\alpha_{n,H_{n}})$. Thus, it holds that:
@@
$\quad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]$
$\quad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})\}(\mathcal{T}_{0}(q_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(q_{n,H_{n}}))\right]$

### v4

@@ H.3 Derivation of von Mises expansion in Theorem 1
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
Using that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $r_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$, we can show that
$E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
@@
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mathcal{T}_{P}(q_{P})(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
The first result then follows. In the case where $P,\widehat{P}\in\mathcal{P}_{H}$, we have that $r_{P}=\mathcal{T}_{P}(q_{P,H})$ and $\widehat{r}=\widehat{\mathcal{T}}(\widehat{q}_{H})$. In this case, the expression simplifies to:
@@
Proof of Theorem 3.
We use the standard fact that the EIF in a restricted model is obtained by projecting the unrestricted (nonparametric) EIF onto the tangent space of that restricted model (bickel1993efficient). By kallus2022efficiently, the nonparametric EIF is
$\varphi_{0}(S_{1},Y_{0},A_{0},S_{0})=d_{0}(A_{0},S_{0})\,\Big\{Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})-q_{0}(A_{0},S_{0})\Big\}+V^{\pi}(q_{0})(S_{0})-\psi_{0}.$
When the law of $S_{0}$ is known, the EIF is obtained by projecting $\varphi_{0}$ onto functions with mean zero given $S_{0}$, yielding
$\widetilde{\varphi}_{0}(S_{1},Y_{0},A_{0},S_{0}):=d_{0}(A_{0},S_{0})\,\Big\{Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})-q_{0}(A_{0},S_{0})\Big\}.$
We may decompose $\widetilde{\varphi}_{0}$ as
$\widetilde{\varphi}_{0}(S_{1},Y_{0},A_{0},S_{0})$ $=\varphi_{0,1}(S_{1},A_{0},S_{0})+\varphi_{0,2}(Y_{0},A_{0},S_{0}),$
where $\varphi_{0,1}$ is mean zero in $(S_{1},A_{0},S_{0})$ given $(A_{0},S_{0})$, and $\varphi_{0,2}$ is mean zero in $(Y_{0},A_{0},S_{0})$ given $(A_{0},S_{0})$. Moreover, $\varphi_{0,1}$ and $\varphi_{0,2}$ are orthogonal in $L^{2}(P_{0})$.
The tangent space for the model where the law of $S_{0}$ and $r_{0}$ are known is the Hilbert space of square-integrable, mean-zero functions of $(S_{1},A_{0},S_{0})$ given $(A_{0},S_{0})$. The projection of $\widetilde{\varphi}_{0}$ onto this space is $\varphi_{0,1}$, and hence $\varphi_{0,1}$ is the EIF in this model.
Similarly, when the law of $S_{0}$ and the transition kernel are known, the tangent space is the Hilbert space of square-integrable, mean-zero functions of $(Y_{0},A_{0},S_{0})$ given $(A_{0},S_{0})$. The projection of $\widetilde{\varphi}_{0}$ onto this space is $\varphi_{0,2}$, so $\varphi_{0,2}$ is the EIF in this model. ∎
Appendix I Proofs for Section 3 on semiparametric DRL
I.1 Asymptotic linearity of DRL estimator under correct specification
The following theorem establishes the pathwise differentiability of the parameter $\Psi$ under the semiparametric model $\mathcal{P}_{H}$. For each $P\in\mathcal{P}_{H}$, we define the influence function
Theorem 12 (Pathwise differentiability).
In the following conditions and theorem, let $\varphi_{n,H}$ denote the estimator of the IF $\varphi_{0,H}$ from Theorem 12, obtained by plugging in our nuisance estimators.
@@ I.1 Asymptotic linearity of DRL estimator under correct specification
I.2 Asymptotic linearity and efficiency of model-robust DRL estimator
@@
$R_{n,H}^{*}(P_{0})$ $:=P_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{n,H})\right\}(r_{0}-r_{n})\right]$
$\qquad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\}(\mathcal{T}_{0}(q_{0,H})-\mathcal{T}_{0}(q_{n,H}^{\pi}))\right]$
$\qquad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\}(\mathcal{T}_{0}(q_{n,H}^{\pi})-\widehat{\mathcal{T}}_{n}(q_{n,H}^{\pi}))\right].$
@@
Appendix J Proofs for Section 4 on calibrated FQI
J.1 Lemma bounding approximation error of estimated features
In the following lemma, let $X$ be a covariate and $Y\in\mathbb{R}$ be an outcome. For feature transformations $\varphi_{n},\varphi_{0}$, denote $f_{(\varphi_{n},\varphi_{0})}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x),\varphi_{0}(X)=\varphi_{0}(x)]$, $f_{\varphi_{n}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x)]$, and $f_{\varphi_{0}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{0}(X)=\varphi_{0}(x)]$. The following is a generalization of Lemma 1 in vanderlaan2024automatic.
@@
$\left|g(\varphi_{n}(x),\varphi_{0}(x))-E[g(\varphi_{n}(X),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$
$=\left|E[g(\varphi_{n}(x),\varphi_{0}(x))-g(\varphi_{n}(x),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$
@@
Now, for some function $f$, suppose that $(\varphi_{n}(x),\varphi_{0}(x))\mapsto(\Pi_{\varphi_{n},\varphi_{0}}f)(x)$ is Lipschitz continuous with fixed constant $L$. Then, defining $g:(\widehat{m},m)\mapsto E_{0}[f(X)|\varphi_{n}(X)=\widehat{m},\varphi_{0}(X)=m,\mathcal{D}_{n}]$ and noting by the law of iterated expectation that $\Pi_{\varphi_{n}}\Pi_{\varphi_{n},\varphi_{0}}f=\Pi_{\varphi_{n}}f$, we obtain the following pointwise error bound:
@@
J.2 Proofs of main results
@@
We prove the theorem for any given $q_{n}$ (e.g., $q_{n}^{*}$) as calibration is not necessary. We introduce the following notation. For any feature transformation $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, let $d_{0,\phi}=\mathcal{T}_{0,\phi}(\alpha_{0,\phi})$, where $\alpha_{0,\phi}:=\operatorname*{argmin}_{\alpha\in\overline{H}_{\phi}}E_{0}\left[\{\mathcal{T}_{0,\phi}(f\circ\phi)(A_{0},S_{0})\}^{2}-2m(S_{0},A_{0},\alpha)\right]$ is the Riesz representer for the function class $H_{\phi}:=\{f\circ\phi;f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$ induced by $\phi$. Denote $r_{0,\phi}$ by the map $(a,s)\mapsto E_{0}[Y_{0}\mid\phi(A_{0},S_{0})=\phi(a,s)]$. Let $\alpha_{0,q_{n}}:=\operatorname*{argmin}_{\alpha\in\overline{H}_{q_{n}}}\left\|\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(\alpha)\right\|_{P_{0}}$. Then, we have $d_{0,q_{n}}=\mathcal{T}_{0,q_{n}}(\alpha_{0,q_{n}})$.
@@
$=\langle d_{0,(q_{n},q_{0})}-d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$\qquad+\langle d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
Note $\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})=r_{0}$ by correct specification of $H_{(q_{n},q_{0})}$ for $q_{0}$. Thus, by the law of total expectation, we have
$\langle d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$ $=\langle d_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$=\langle d_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-r_{0}\rangle_{P_{0}}$
@@
where the final equality follows from the orthogonality conditions of the Bellman projection $q_{0,q_{n}}$ of $q_{0}$ with respect to the norm $\|\mathcal{T}_{0,q_{n}}(\cdot)\|_{P_{0}}$ and the fact that $d_{0,q_{n}}$ is in the closed range of $\mathcal{T}_{0,q_{n}}$. Hence,
$\Psi_{n}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n},q_{0})}-d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
@@
$|\Psi_{n}(P_{0})-\Psi(P_{0})|$ $\leq\|d_{0,(q_{n},q_{0})}-d_{0,q_{n}}\|_{P_{0}}\cdot\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}}.$
For the second part of the theorem, take $d_{0,q_{n}}$ equal to $\mathcal{T}_{0,q_{n}}(\alpha_{0,q_{n}})$. By the invertibility condition in D1, $\mathcal{T}_{0,q_{n}}(\overline{H}_{q_{n}})$ equals $\overline{H}_{q_{n}}$ from the first part of (D2). Hence, we have that $d_{0,q_{n}}$ equals the projection $\operatorname*{argmin}_{d\in\overline{H}_{q_{n}}}\|d_{0,(q_{n},q_{0})}-d\|_{P_{0}}.$ Next, using the second part of (D2), we apply Lemma 14 with $Y:=d_{0,(q_{n},q_{0})}$, $\varphi_{n}:=q_{n}$, and $\varphi_{0}:=q_{0}$ to conclude that
$\|d_{0,(q_{n},q_{0})}-d_{0,q_{n}}\|_{P_{0}}\lesssim\|q_{n}-q_{0}\|_{P_{0}}.$
@@
Lemma 15 (Calibration corrects plug-in bias).
Suppose $q_{n}^{*}$ is Bellman calibrated (10). Then,
@@
Taking $f$ such that $f\circ q_{n}^{*}=d_{0,q_{n}^{*}}$, we find that
$\sum_{i=1}^{n}d_{0,q_{n}^{*}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
Therefore, the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*})$ is equal to the DRL estimator:
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*})+\sum_{i=1}^{n}d_{0,q_{n}^{*}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}.$
@@
$\qquad+\int\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma V^{\pi}(q_{n}^{*})(a,s^{\prime})-q_{n}^{*}(a,s)\}dP_{0}(s,a,y,s^{\prime})$
@@
$\qquad+\int\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma V^{\pi}(q_{n}^{*})(a,s^{\prime})-q_{n}^{*}(a,s)\}dP_{0}(s,a,y,s^{\prime}),$
@@
$\int\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma V^{\pi}(q_{n}^{*})(a,s^{\prime})-q_{n}^{*}(a,s)\}dP_{0}(s,a,y,s^{\prime})$ $=\langle\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}}),\mathcal{T}_{0,q_{n}^{*}}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,q_{n}^{*}}(q_{n}^{*})\rangle_{P_{0}}.$
@@
$\qquad+\langle{d}_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{n}^{*})\rangle_{P_{0}}$
@@
Appendix K Proofs for ADML in Section F
@@
Observe that $\psi_{n,H_{n}}=\Psi_{H_{n}}(\widehat{P_{n}})+P_{n}\varphi_{n,H_{n}}$, where $\widehat{P}_{n}\in\mathcal{P}$ is any distribution such that $q_{\widehat{P}_{n},H_{n}}=q_{n,H_{n}}$, $r_{\widehat{P}_{n}}=\widehat{\mathcal{T}}_{n}(q_{n,H_{n}})$, and $T_{\widehat{P}_{n}}(\alpha_{\widehat{P}_{n},H_{n}})=T_{\widehat{P}_{n}}(\alpha_{n,H_{n}})$. Thus, it holds that:
@@
$\qquad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]$
$\qquad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})\}(\mathcal{T}_{0}(q_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(q_{n,H_{n}}))\right]$